Circular No. 8926

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2008 A3 (SOHO) Further to IAUC 8923, an additional tiny, very faint (mag ∼ 7.5) comet

was found on SOHO website images that is a member of the Marsden group (“discovery” observation tabulated below); according to K. Battams, it may have been “borderline diffuse”, and it maintained a very constant brightness level until quite near the end, when it started to fade. R. Kracht suggested the identity of C/2008 A3 with C/2002 R1 (cf. IAUC 7969, MPEC 2002-R57), and B. G. Marsden published a linked orbit on MPEC 2008-B61 that yields T = 1997 Apr. 19.0 for its previous (unobserved) perihelion passage.

Comet 2008 UT α_{2000} δ_{2000} Inst. F MPEC
C/2008 A3 Jan. 15.588 19^{h}52.^{m}2 −20^{o}39′ C2 RK 2008-B61

COMETS C/2008 D1, C/2008 D2, C/2008 D3, C/2008 D4 (STEREO)

K. Battams and K. Baldwin have reported measurements, reduced and analyzed by B. G. Marsden, for four Kreutz-sungrazing comets found on images obtained with NASA’s new “Solar-Terrestrial Relations Observatory” (STEREO), which involves two satellites ultimately to be located 45^{o} ahead (STEREO-A) and 45^{o} behind (STEREO-B) the earth in its orbit. The SECCHI suite of instruments on each satellite contains two coronagraphs (‘COR2’ having bandpass 650–750 nm and resolution 15′′/pixel) and a “Heliospheric Imager” consisting of two telescopes (‘HI1’ having bandpass 650–750 nm and resolution 35′′/pixel). The tabulated “discovery” observations below are given in the same format as used for the SOHO comets (see above); the ‘C’ under instrumentation denotes the ‘COR2’ instrument on both STEREO-A and -B, while the ‘HI’ flag denotes the ‘HI1-A’ instrument only. C/2008 D1, C/2008 D2, and C/2008 D4 were very faint (mag ∼ 10–11 for the first and third objects, and ∼ 9 for the second) and diffuse. C/2008 D3 was of mag perhaps 7 with a thin, faint tail in HI1-A images, and brighter (mag perhaps 3) with a thin tail in COR2 images; this is the first near-sun comet to have an orbit determined (cf. MPEC 2008-E58) using astrometry from two well-separated spacecraft (now 0.76 AU apart).

Comet 2008 UT α_{2000} δ_{2000} Inst. F MPEC
C/2008 D1 Feb. 16.822 23^{h}43.^{m}4 − 5^{o}01′ HI  AW 2008-E58
C/2008 D2 19.461 23 55.6 − 4 05 HI AW 2008-E58
C/2008 D3  20.294 0 14.0 − 4 49 HI/C AW 2008-E58
C/2008 D4  20.933 23 58.9 − 3 23 HI  RK 2008-E58
2008 March 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green